# Spontaneous Magnon Decays from Nonrelativistic Time-Reversal Symmetry Breaking in Altermagnets

### v1

@@
$\mathcal{D}_{\pm}(\omega,{\boldsymbol{k}})=\frac{1}{N_{\textrm{muc}}^{2}}\sum_{{\boldsymbol{p}},{\boldsymbol{q}}}\delta\left(\omega-\epsilon_{\pm}({\boldsymbol{p}})-\epsilon_{\pm}({\boldsymbol{q}})-\epsilon_{\mp}({\boldsymbol{k}}-{\boldsymbol{p}}-{\boldsymbol{q}})\right)$ (10)
is the density of states of the three-magnon continuum built from three magnons whose spin adds up to ±$\pm 1$. Carrying a finite net spin, the $\mathcal{D}_{\pm}(\omega,{\boldsymbol{k}})$ continua also exhibit the d-wave splitting. Whenever these continua overlap with the single-magnon branch of the same spin, that is, whenever $\mathcal{D}_{\pm}(\epsilon_{\pm}({\boldsymbol{k}}),{\boldsymbol{k}})\neq 0$, spontaneous decays are kinematically possible [3].
@@
Figure 4a shows $\mathcal{D}_{+}(\omega,{\boldsymbol{k}})$ together with the noninteracting magnon spectrum for $\kappa<0$ and $\Delta=0.3$. The spin-$\uparrow$ magnon with energy $\epsilon_{+}({\boldsymbol{k}})$ is embedded within the $\mathcal{D}_{+}(\omega,{\boldsymbol{k}})$ continuum along the direction $(0,0)$ to $(\frac{\pi}{2},-\frac{\pi}{2})$, and can therefore decay. In contrast, along the direction $(0,0)$ to $(\frac{\pi}{2},\frac{\pi}{2})$ the continuum begins exactly at that branch’s energy (due to the Goldstone mode), precluding decays. A similar discussion applies to the spin-$\downarrow$ magnon with energy $\epsilon_{-}({\boldsymbol{k}})$ and the $\mathcal{D}_{-}(\omega,{\boldsymbol{k}})$ continuum (not shown). Along the nodal-plane direction from $(0,0)$ to $(\pi,0)$, the continuum also begins exactly at the single-particle energies because $\Delta<\Delta^{\star}_{\text{d-wave}}$, rendering magnons stable.
@@
For $\kappa<0$ and $\Delta=1.0$, the lower magnon branch exhibits a local minimum at $(-\frac{\pi}{2},\frac{\pi}{2})$, which causes a rapid increase in the three-particle DOS at about three times its energy, as shown in Fig. 4b. As a result, $\Gamma({\boldsymbol{k}})$ of the upper branch also exhibits a sharp increase at this energy, in agreement with what is seen in Fig. 3j. Since $\Delta>\Delta^{\star}_{\text{d-wave}}$, the magnon energies are embedded within the continua along the nodal-plane direction from $(0,0)$ to $(\pi,0)$, explaining the finite decay rate seen in Fig. 3j.
@@
For $\kappa>0$, kinematics predicts an island of stability of size $k^{\star}_{\text{d-wave}}$ for the lower magnon branch around the origin. Indeed, as shown in Fig. 4c, the relevant continua for the lower magnon mode exhibit a pronounced drop to zero with d-wave-shaped lobes. Around the origin, the density of states of the continuum stays finite only along the nodal planes, where the altermagnetic splitting cannot protect the lower branch from decays.
@@
The existence of Goldstone modes implies universal low-energy physics, raising the question how $\Gamma_{\pm}({\boldsymbol{k}})$ in Eq. (9) scales with $|{\boldsymbol{k}}|$ as $|{\boldsymbol{k}}|\to 0$. Its scaling depends on both that of the available decay phase space, i.e., the scaling of the on-shell three-magnon density of states $\mathcal{D}_{\pm}(\epsilon_{\pm}({\boldsymbol{k}}),{\boldsymbol{k}})$—as given by Eq. (10) with $\omega\to\epsilon_{\pm}({\boldsymbol{k}})$—and that of $W_{\pm}({\boldsymbol{k}},{\boldsymbol{q}},{\boldsymbol{p}})$.
@@
We begin with the decay phase space, which we call $V({\boldsymbol{k}})$ for brevity. At long wavelengths, the linear magnon spectrum $\epsilon({\boldsymbol{k}})=v|{\boldsymbol{k}}|$ is the reference point. A magnon at ${\boldsymbol{k}}$ is “ready to decay” into magnons at ${\boldsymbol{p}}$, ${\boldsymbol{q}}$, and ${\boldsymbol{k}}-{\boldsymbol{p}}-{\boldsymbol{q}}$ along the line from the origin to ${\boldsymbol{k}}$. Hence, any positive curvature—be it due to a quadratic or cubic nonlinearity—activates decays of the type in Fig. 2c in a threshold behavior. Since the two momenta ${\boldsymbol{q}}$ and ${\boldsymbol{p}}$ have to be almost aligned with ${\boldsymbol{k}}$ to fulfill energy and momentum conservation, the four-dimensional decay phase space has a cigar shape. In the two long directions, it is constrained by $|{\boldsymbol{k}}|$ and in the other two short directions by $|{\boldsymbol{k}}|^{3/2}$. It follows that $V({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{3}$ for a positively curved magnon dispersion. In case of altermagnets with their even-parity spin split dispersion $\epsilon_{\pm}({\boldsymbol{k}})=v|{\boldsymbol{k}}|\pm\Delta\sin(n\varphi_{{\boldsymbol{k}}})|{\boldsymbol{k}}|^{n}$, this observation implies that $V({\boldsymbol{k}})$ of the upper branch is zero at $\Delta=0$ but jumps to a finite value at $\Delta\neq 0$.
@@
The threshold value of $\Delta$ becomes nonzero and $n$-dependent once additional nonlinearities are taken into account. Including a negative cubic nonlinearity, $\kappa<0$, as in Eq. (3), we find that $\Delta$ has to exceed $-C_{n}\kappa|{\boldsymbol{k}}|^{3-n}$, with $C_{n}>0$ being a constant, to cause decays of the upper branch at $|{\boldsymbol{k}}|$. For d-wave magnets ($n=2$), we find that $\Delta>-C_{n}\kappa|{\boldsymbol{k}}|$ holds for any value of $\kappa<0$ if only $|{\boldsymbol{k}}|$ is small enough. Thus, upper-branch magnons in d-wave magnets have a decay phase space $V_{\text{d-wave}}({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{3}$ for $\kappa<0$ as $|{\boldsymbol{k}}|\to 0$. In contrast, g-wave and i-wave magnets have a negative power of $|{\boldsymbol{k}}|^{3-n}$, implying that their threshold value for $\Delta$ diverges as $|{\boldsymbol{k}}|\to 0$. Thus, the upper magnon branch of g-wave and i-wave magnets with $\kappa<0$ does not exhibit the $V({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{3}$ long-wavelength scaling. Instead, other higher-order-in-momentum mechanisms have to be identified; an example is the mechanism depicted in Fig. 2d.
@@
With the relevant decay process for d-wave altermagnets identified, we plug the “cigar-shaped” scaling ($\propto|{\boldsymbol{k}}|$ for the two long and $\propto|\Delta|^{1/2}|{\boldsymbol{k}}|^{3/2}$ for the two short directions) for the respective components of ${\boldsymbol{p}}$ and ${\boldsymbol{q}}$ into $W_{\pm}({\boldsymbol{k}},{\boldsymbol{q}},{\boldsymbol{p}})$ and take the limit $|{\boldsymbol{k}}|\to 0$. We find $W_{\pm}({\boldsymbol{k}},{\boldsymbol{q}},{\boldsymbol{p}})\propto|\Delta||{\boldsymbol{k}}||\sin(2\varphi_{\boldsymbol{k}})|$, where we have restored the angular dependence. Therefore, by combining the square of the vertex with the decay phase space scaling according to Eq. (9), the damping of the upper magnon branch in d-wave altermagnets is given by
@@
$\Gamma_{\text{upper}}({\boldsymbol{k}})\propto|\Delta|^{2}|{\boldsymbol{k}}|^{5}|\sin(2\varphi_{\boldsymbol{k}})|^{2}\quad\text{as}\quad|{\boldsymbol{k}}|\to 0,$ (11)
as indicated in Fig. 4d.
@@
Magnons along the nodal planes are stable if $\Delta<\Delta^{\star}_{\text{d-wave}}$ but exhibit the decays shown in Fig. 2a for $\Delta>\Delta^{\star}_{\text{d-wave}}$, with $\Delta^{\star}_{\text{d-wave}}$ given in Eq. (5). In this case, the cigar-shaped decay space has a short-axes scaling of $|{\boldsymbol{k}}|^{2}$, such that the damping along the nodes is given by
@@
which is also indicated in Fig. 4d. $\Theta$ indicates the Heaviside step function. Also, for $\kappa>0$, an isotropic scaling $\Gamma({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{7}$ is expected [11].
We conclude that in the case of d-wave altermagnets, the cubic nonlinearity is subleading in the long-wavelength limit and away from the nodal planes, giving rise to the unique leading $|{\boldsymbol{k}}|^{5}$ scaling of the decay rate in Eq. (11). For further details on the long-wavelength scaling of magnon damping, see Supplementary Notes 3 and 4. We emphasize that our results hold for two-dimensional altermagnets.
@@
Figure 5: Comparison of nonperturbative and perturbative transverse quantum dynamics in two-dimensional d-wave altermagnets. (a) The transverse dynamical structure factor $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$, as obtained from DMRG$+$tMPO, shows clear altermagnetic splitting along the path $(0,0)$ to $(\frac{\pi}{2},\frac{\pi}{2})$ in reciprocal space. Around this point, the spectral weight of the upper magnon branch gets almost completely wiped out due to hybridization with continua. A small XXZ anisotropy $\lambda=0.02$ was added for numerical stability. It causes a tiny gap of the Goldstone modes. Note the double linear color bar scale. Dashed white lines indicate points at which simulation results for different simulation geometries were patched together (see Methods). As the simulation on cylinders weakly breaks the exact four-fold rotational symmetry of the lattice, small discontinuities in the intensity can be identified. (b) The single-particle spectral function $\mathcal{A}({\boldsymbol{k}},\omega)$ as obtained within perturbative nonlinear spin-wave (NLSW) theory of $1/S^{2}$ accuracy modulo additional off-shell corrections. The colors on the horizontal momentum axes in (a) and (b) match the colors of the high-symmetry paths indicated in the Brillouin zone in the inset of (b). (c,d) Line cuts of the nonperturbative $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ versus frequency at momenta $(\frac{\pi}{2},\frac{\pi}{2})$ and $(\pi,0)$, respectively. (e,f) Line cuts of the perturbative $\mathcal{A}({\boldsymbol{k}},\omega)$ versus frequency at momenta $(\frac{\pi}{2},\frac{\pi}{2})$ and $(\pi,0)$, respectively. Parameters read $J_{1}=1$, $J_{2}=-0.5$, $S=1/2$, and $\Delta$ as indicated.
@@
The observation of magnon decays arising from the nonrelativistic altermagnetic beyond-s-wave band splitting was derived within the framework of perturbation theory. At low energies, where both the decay phase space and magnon-magnon interaction vertices are small, perturbation theory provides reliable results [8, 9]. At higher energies, our calculations suggest $\Gamma/\epsilon\lesssim 0.1$ for $S=1/2$, indicating that magnons remain well-defined, albeit damped, quasiparticles, validating the perturbative approach. However, small spin $S$, non-universal high-energy decays, and enhanced quantum fluctuations in low-dimensional systems necessitate nonperturbative methods to fully assess magnon stability. We address this using density matrix renormalization group (DMRG) [26, 27] and matrix product operator time evolution (tMPO) [28] techniques to extract the dynamical spin structure factor $\mathcal{S}_{\mu\nu}({\boldsymbol{k}},\omega)$ (with $\mu$ and $\nu$ denoting $x$, $y$, and $z$), as measured in inelastic neutron scattering (see Methods for details).
@@
Figure 5a presents the transverse dynamical structure factor $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)=\mathcal{S}_{xx}({\boldsymbol{k}},\omega)+\mathcal{S}_{yy}({\boldsymbol{k}},\omega)$ from nonperturbative simulations, revealing distinct magnon peaks. Peak intensity decreases due to magnon damping and non-Bravais lattice form factors, explaining the vanishing intensity at $(0,0)$. The lower mode remains sharp (within Gaussian filtering of width $\sigma_{\omega}=0.0896$), while the upper mode broadens along $(0,0)\to(\frac{\pi}{2},\frac{\pi}{2})$, aligning with perturbation theory. At $(\frac{\pi}{2},\frac{\pi}{2})$, its non-Lorentzian shape at large $\Delta$ signals strong nonperturbative effects.
@@
We find consistently good agreement between the nonperturbative spectrum in Fig. 5a and the $1/S^{2}$ off-shell single-particle spectral function $\mathcal{A}({\boldsymbol{k}},\omega)$ in Fig. 5b (see Methods). Linecuts through $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ and $\mathcal{A}({\boldsymbol{k}},\omega)$ in Fig. 5c,d and Fig. 5e,f at $(\frac{\pi}{2},\frac{\pi}{2})$ and $(\pi,0)$ show remarkable agreement. However, perturbation theory underestimates nonperturbative high-energy tails (compare Fig. 5d with Fig. 5f), possibly due to three-particle contributions, spinon continua [29], or Higgs resonances [30]. Capturing the full quantum dynamics of $S=1/2$ altermagnets may thus require incorporating amplitude oscillations of the order parameter, which lie beyond perturbation theory.

### v2

@@ Results
$\mathcal{D}_{\pm}(\omega,{\boldsymbol{k}})=\frac{1}{N_{\textrm{muc}}^{2}}\sum_{{\boldsymbol{p}},{\boldsymbol{q}}}\delta\left(\omega-\epsilon_{\pm}({\boldsymbol{p}})-\epsilon_{\pm}({\boldsymbol{q}})-\epsilon_{\mp}({\boldsymbol{k}}-{\boldsymbol{p}}-{\boldsymbol{q}})\right).$ (10)
@@
When these continua overlap with the single-magnon branch of the same spin, that is, when $\mathcal{D}_{\pm}(\epsilon_{\pm}({\boldsymbol{k}}),{\boldsymbol{k}})\neq 0$, spontaneous decays are kinematically possible [3]. Figure 4a shows $\mathcal{D}_{+}(\omega,{\boldsymbol{k}})$ together with $\epsilon_{\pm}({\boldsymbol{k}})$ for $\kappa<0$ and $\Delta=0.3$. Clearly, the spin-$\uparrow$ magnon $\epsilon_{+}({\boldsymbol{k}})$ is embedded within $\mathcal{D}_{+}(\omega,{\boldsymbol{k}})$ along $(0,0)$ to $(\frac{\pi}{2},-\frac{\pi}{2})$, allowing decays. In contrast, along $(0,0)$ to $(\frac{\pi}{2},\frac{\pi}{2})$ the continuum begins exactly at that branch’s energy (due to the Goldstone mode), precluding decays. A similar discussion applies to the spin-$\downarrow$ magnon $\epsilon_{-}({\boldsymbol{k}})$ and $\mathcal{D}_{-}(\omega,{\boldsymbol{k}})$ (not shown). Along the nodal-plane from $(0,0)$ to $(\pi,0)$, the continuum also begins exactly at the single-particle energies because $\Delta<\Delta^{\star}_{\text{d-wave}}$, ensuring stability.
@@
For $\kappa<0$ and $\Delta=1.0$, the lower magnon branch has a local minimum at $(\pm\frac{\pi}{2},\frac{\pi}{2})$, causing a rapid increase in the three-particle density of states at approximately three times its energy, as shown in Fig. 4b. This results in a sharp increase in $\Gamma({\boldsymbol{k}})$ for the upper branch, in agreement with Fig. 3j. Since $\Delta>\Delta^{\star}_{\text{d-wave}}$, magnon energies are embedded within the continua along the nodal planes from $(0,0)$ to $(\pi,0)$, explaining the finite decay rate in Fig. 3j.
@@
For $\kappa>0$, kinematics predicts an island of stability of size $k^{\star}_{\text{d-wave}}$ for the lower magnon branch around the origin. As shown in Fig. 4c, the relevant continua exhibit a drop to zero with d-wave-shaped lobes. Near the origin, the density of states remains finite only along the nodal planes, where the altermagnetic splitting does not protect the lower branch from decay.
@@
The presence of Goldstone modes raises the question of how $\Gamma_{\pm}({\boldsymbol{k}})$ in Eq. (9) scales with $|{\boldsymbol{k}}|$ as $|{\boldsymbol{k}}|\to 0$. This scaling depends on the decay phase space (denoted by $V({\boldsymbol{k}})$), specifically the scaling of the three-magnon density of states $\mathcal{D}_{\pm}(\epsilon_{\pm}({\boldsymbol{k}}),{\boldsymbol{k}})$, and that of $W_{\pm}({\boldsymbol{k}},{\boldsymbol{q}},{\boldsymbol{p}})$.
@@
At long wavelengths, the linear magnon spectrum $\epsilon({\boldsymbol{k}})=v|{\boldsymbol{k}}|$ is the reference point. A magnon at $\boldsymbol{k}$} is “ready to decay” into magnons at $\boldsymbol{p}$}, $\boldsymbol{q}$}, and ${\boldsymbol{k}}-{\boldsymbol{p}}-{\boldsymbol{q}}$, with any positive curvature activating decays in a threshold behavior, as in Fig. 2c. Since the two momenta ${\boldsymbol{p}}=(p_{x},p_{y})$ and ${\boldsymbol{q}}=(q_{x},q_{y})$ have to be almost aligned with $\boldsymbol{k}$} to fulfill energy and momentum conservation, the four-dimensional decay phase space spanned by $\{p_{x},p_{y},q_{x},q_{y}\}$ has a cigar shape. In the two long directions, it is constrained by $|{\boldsymbol{k}}|$, and in the other two short directions by $|{\boldsymbol{k}}|^{3/2}$, yielding $V({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{3}$ for a positively curved magnon dispersion. In altermagnets with magnon dispersion $\epsilon_{\pm}({\boldsymbol{k}})=v|{\boldsymbol{k}}|\pm\Delta\sin(n\varphi_{{\boldsymbol{k}}})|{\boldsymbol{k}}|^{n}$, this implies $V({\boldsymbol{k}})$ of the upper branch is zero at $\Delta=0$ and jumps to a finite value at $\Delta\neq 0$.
@@
The threshold value of $\Delta$ becomes nonzero and $n$-dependent once additional nonlinearities are taken into account. When a negative cubic nonlinearity ($\kappa<0$) is included as in Eq. (3), $\Delta$ has to exceed $C_{n}\kappa|{\boldsymbol{k}}|^{3-n}$ (with $C_{n}>0$ being a constant) for the upper branch to decay at $|{\boldsymbol{k}}|$. For d-wave magnets ($n=2$), $\Delta>-C_{n}\kappa|{\boldsymbol{k}}|$ for any $\kappa<0$ as long as $|{\boldsymbol{k}}|$ is small enough and their upper-branch magnons have a decay phase space $V_{\text{d-wave}}({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{3}$. In contrast, the threshold value for $\Delta$ diverges as $|{\boldsymbol{k}}|\to 0$ in g-wave and i-wave magnets, implying that their upper magnon branch does not exhibit the $V({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{3}$ long-wavelength scaling for $\kappa<0$. Instead, other higher-order-in-momentum mechanisms, e.g., the one in Fig. 2d, have to be identified.
@@
With the relevant decay process for d-wave altermagnets identified, we plug the “cigar-shaped” scaling ($\propto|{\boldsymbol{k}}|$ for the two long and $\propto|\Delta|^{1/2}|{\boldsymbol{k}}|^{3/2}$ for the two short directions) for the respective components of $\boldsymbol{p}$} and $\boldsymbol{q}$} into $W_{\pm}({\boldsymbol{k}},{\boldsymbol{q}},{\boldsymbol{p}})$ and take $|{\boldsymbol{k}}|\to 0$. We find $W_{\pm}({\boldsymbol{k}},{\boldsymbol{q}},{\boldsymbol{p}})\propto|\Delta||{\boldsymbol{k}}||\sin(2\varphi_{\boldsymbol{k}})|$, where we have restored the angular dependence. Therefore, as indicated in Fig. 4d, the damping of the upper magnon branch is
@@
$\Gamma_{\text{upper}}({\boldsymbol{k}})\propto|\Delta|^{2}|{\boldsymbol{k}}|^{5}|\sin(2\varphi_{\boldsymbol{k}})|^{2}\quad\text{as}\quad|{\boldsymbol{k}}|\to 0.$ (11)
@@
Magnons along the nodal planes are stable if $\Delta<\Delta^{\star}_{\text{d-wave}}$ but exhibit the decays shown in Fig. 2a for $\Delta>\Delta^{\star}_{\text{d-wave}}$, with $\Delta^{\star}_{\text{d-wave}}$ given in Eq. (5). In this case, the cigar-shaped decay space has a short-axes scaling of $|{\boldsymbol{k}}|^{2}$, such that the damping along the nodes is
@@
where $\Theta$ indicates the Heaviside step function. Also, for $\kappa>0$, an isotropic scaling $\Gamma({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{7}$ is expected [11].
Therefore, in d-wave altermagnets, the cubic nonlinearity is subleading at long-wavelengths and away from the nodal planes, giving rise to a unique scaling $\Gamma\propto|{\boldsymbol{k}}|^{5}$ in two dimensions. For further details, see Supplementary Material III and IV.
@@
Figure 5: Comparison of nonperturbative and perturbative transverse quantum dynamics in two-dimensional d-wave altermagnets. (a) The transverse dynamical structure factor $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$, as obtained from DMRG++tMPO, shows clear altermagnetic splitting along the path $(0,0)$ to $(\frac{\pi}{2},\frac{\pi}{2})$ in reciprocal space. Around this point, the spectral weight of the upper magnon branch gets almost completely wiped out due to hybridization with continua. A small XXZ anisotropy $\lambda=0.02$ was added for numerical stability. It causes a tiny gap of the Goldstone modes. Note the double linear color bar scale. Dashed white lines indicate points at which simulation results for different simulation geometries were patched together (see Appendix C). As the simulation on cylinders weakly breaks the exact four-fold rotational symmetry of the lattice, small discontinuities in the intensity can be identified. (b) The single-particle spectral function $\mathcal{A}({\boldsymbol{k}},\omega)$ as obtained within perturbative nonlinear spin-wave (NLSW) theory of $1/S^{2}$ accuracy modulo additional off-shell corrections. The colors on the horizontal momentum axes in (a) and (b) match the colors of the high-symmetry paths indicated in the Brillouin zone in the inset of (b). (c,d) Line cuts of the nonperturbative $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ versus frequency at momenta $(\frac{\pi}{2},\frac{\pi}{2})$ and $(\pi,0)$, respectively. (e,f) Line cuts of the perturbative $\mathcal{A}({\boldsymbol{k}},\omega)$ versus frequency at momenta $(\frac{\pi}{2},\frac{\pi}{2})$ and $(\pi,0)$, respectively. Parameters read $J_{1}=1$, $J_{2}=-0.5$, $S=1/2$, and $\Delta$ as indicated.
@@
At low energies, where decay phase space and magnon-magnon interaction vertices are small, perturbation theory is reliable [8, 9]. At higher energies, our calculations give $\Gamma/\epsilon\lesssim 0.1$ for $S=1/2$, indicating that magnons remain well-defined quasiparticles, validating the perturbative approach. However, small $S$, non-universal high-energy decays, and enhanced fluctuations in low-dimensional systems require nonperturbative methods for assessing magnon stability. We use density matrix renormalization group (DMRG) [28, 29] and matrix product operator time evolution (tMPO) [30] to compute the dynamical spin structure factor $\mathcal{S}_{\mu\nu}({\boldsymbol{k}},\omega)$, as measured in inelastic neutron scattering (see Appendix C).
@@
Figure 5a shows the transverse $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)=\mathcal{S}_{xx}({\boldsymbol{k}},\omega)+\mathcal{S}_{yy}({\boldsymbol{k}},\omega)$ from nonperturbative simulations, revealing distinct magnon peaks. Vanishing spectral weight at $(0,0)$ is attributed to the disappearance of non-Bravais lattice form factors in the long-wavelength limit. While peak intensity at finite momenta is, in principle, finite due to nonzero form factors, it decreases significantly under the influence of magnon damping. The lower mode remains sharp (within Gaussian filtering of width $\sigma_{\omega}=0.0896$), while the upper mode broadens along $(0,0)\to(\frac{\pi}{2},\frac{\pi}{2})$, aligning with perturbation theory. At $(\frac{\pi}{2},\frac{\pi}{2})$, its non-Lorentzian shape at large $\Delta$ signals strong nonperturbative effects.
@@
We find good agreement between the nonperturbative spectrum in Fig. 5a and the $1/S^{2}$ off-shell single-particle spectral function $\mathcal{A}({\boldsymbol{k}},\omega)$ in Fig. 5b (see Appendix A). Linecuts through $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ and $\mathcal{A}({\boldsymbol{k}},\omega)$ in Fig. 5c,d and Fig. 5e,f at $(\frac{\pi}{2},\frac{\pi}{2})$ and $(\pi,0)$ show remarkable agreement. However, perturbation theory underestimates nonperturbative high-energy tails (compare Fig. 5d with Fig. 5f), possibly due to three-particle contributions, spinon continua [31], or Higgs resonances [32]. Capturing the full quantum dynamics of $S=1/2$ altermagnets may thus require incorporating amplitude oscillations of the order parameter, which lie beyond perturbation theory.
Discussion